OpenStudy (anonymous):

solve by the elimination method 5x-y=41 7y=5x-47

OpenStudy (anonymous):

Are you familiar with the elimination method?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

With the elimination method, you choose a variable to eliminate from your equations and then solve for the other variable. So for example pick one of your two variables to remove.

OpenStudy (anonymous):

ok i choose to get rid of y what do i next

OpenStudy (anonymous):

The x would be easier, but y works fine. So to eliminate the y, you need to add 7 times the first equation to your second equation. 7y = 5x -47 + 7(5x -y) = 7(41) ==============

OpenStudy (anonymous):

Because you can see when you add those two together you will have 0 y's left.

OpenStudy (anonymous):

\[7y + 7(5x -y) = 5x - 47 + 287\] \[\implies 7y + 35x - 7y = 5x +240\] \[\implies 30x = 240\implies x = 8\]